# Nondeterministic Polynomial-time Problem Challenge: An Ever-Scaling Reasoning Benchmark for LLMs

### v1

@@ C.1 Problem Suite: npgym
Interface.
We introduce npgym, a problem suite containing 25 NPC problems with a unified gym-style interface for instance generation and solution verification. Each environment is defined by a problem name and its corresponding hyperparameters, enabling the generation of unlimited problem instances and example solutions. Difficulty can be scaled by adjusting these parameters. npgym also supports automatic verification of solutions produced by large language models (LLMs). New problems can be added easily by implementing two core functions and providing a problem description for prompt generation.
@@
Variables to Scale.
Table 4 lists the variables to scale for each of the 25 NP-complete problems.
@@
Difficulty Levels.
We define and release problem-specific difficulty levels for each of the 25 core problems included in our benchmark. Each problem includes approximately 10 levels of increasing complexity, determined primarily by theoretical factors such as search space size and validated through empirical testing using DeepSeek-R1 and GPT-4o. npgym allows seamless extension to higher difficulty levels as more powerful models become available.
@@
Solution Errors.
There are two fundamental error categories: problem-independent errors and problem-dependent errors. Problem-independent errors are general errors that arise from external factors unrelated to the problem’s intrinsic characteristics and all problems have these types of errors. Problem-independent errors include JSON ERROR (JSON not found or JSON parsing errors), and VERIFICATION ERROR (output format mismatches or structural validation failures). Problem-dependent errors originate from the problem’s inherent complexity, which are defined based on problem specificity. A comprehensive illustration of the errors is displayed in Table 5.
@@ C.2 Solver Suite: npsolver
Online.
The online state-of-the-art LLMs, e.g., o1/o3-mini and DeepSeek-v3/R1, can be accessed through APIs without local computational overhead. However, these online models have dependency on network stability and API costs with token usage. npsolver supports multiple providers, e.g., OpenAI, through modular API clients. We implement efficient batch processing with LiteLLM, which minimizes the latency during parallel problem-solving.
Offline.
Open-weight LLMs, e.g., QwQ-32B and Deepseek-R1-32B, can be accessed by deploying them locally. This allows for GPU-accelerated, high-throughput inference while avoiding API-related costs. Offline models are deployed using vLLM, with hyperparameters—such as temperature and maximum token length—manually configured according to their official technical documentation.
@@ C.3 Evaluation Suite: npeval
npeval employs a statistically rigorous sampling strategy. For each difficulty, the aggregated performance over 3 different independent seeds, with 30 samples generated per seed, aligning with the minimum sample size for reliable statistical analysis [16], are considered. This sampling design, i.e., sampling 90 instances total per difficulty level for each problem, balances budget constraints while mitigating instance-specific variance.
Evaluation Metrics.
rliable [2] is an open-source Python library designed to enable statistically robust evaluation of reinforcement learning and machine learning benchmarks. Inspired by rliable, npeval provides the following 4 evaluation aggregate metrics:
@@
To quantify uncertainty in aggregate metrics, e.g. IQM, npeval employs stratified bootstrap confidence intervals (SBCIs) [11, 12] for the performance interval estimation. SBCIs use stratified resampling within predefined strata, e.g., difficulty levels, to preserve the hierarchical structure of the evaluation data, reduce bias, and provide statistically sound interval estimates.
Comprehensive Analysis
Based on evaluation metrics, npeval provides a comprehensive analysis of the LLMs’ performance over the problems and difficulty levels, including the full results for each problem, each model and each level (Appendix G), the performance over different problems (Appendix H), the analysis of both prompt and completion tokens of LLMs (Appendix I), the analysis of the number of “aha moments” during the DeepSeek-R1 reasoning [14] (Appendix J), an illustration of errors over problems (Table 5) with detailed error analysis (Appendix K), considering both the solution errors, i.e., the errors returned by npgym, and the reasoning errors, i.e., the errors produced in the internal reasoning process of LLMs, which enables the identification of the failure cases (Appendix L). The total cost of evaluation over different models is in listed in Table 23 (Appendix M).
@@ Appendix D Prompts and Responses
The code extracts JSON data from LLM responses using three regex patterns in sequence:
  • First tries to find content between triple quotes with "json" marker,
  • If that fails, looks for "json" followed by content in curly braces,
@@
If all the three tries cannot find the content, we will raise the error.
@@ Appendix E List of NP-complete Problems
  • Input: A set of $m$ clauses $\{C_{1},C_{2},\dots,C_{m}\}$ over a set of $n$ Boolean valued variables $X_{n}=\{x_{1},x_{2},\dots,x_{n}\}$, such that each clause depends on exactly three distinct variables from $X_{n}$. A clause being a Boolean expression of the form $y_{i}\land y_{j}\land y_{k}$ where each $y$ is of the form $x$ or $\neg x$ (i.e. negation of $x$) with $x$ being some variable in $X_{n}$. For example if $n=4$ and $m=3$, a possible instance could be the (set of) Boolean expressions: $C_{1}=(x_{1}\land(\neg x_{2})\land(\neg x_{3}))$, $C_{2}=(x_{2}\land x_{3}\land(\neg x_{4}))$, $C_{3}=((\neg x_{1})\land x_{3}\land x_{4})$.
@@
  • Question: Are there two positive integers $x$ and $y$ such that $(a*x*x)+(b*y)=c$?
@@
  • Question: Is there a linear ordering of $V$ with bandwidth at most $k$, i.e., a one-to-one function $f:V\rightarrow\{0,1,2,...,n-1\}$ such that for all edges ${u,v}$ in $G$, $|f(u)-f(v)|\leq k$?
@@
[…]
  • Question: Is there an ordering $\langle\pi(1),\pi(2),...,\pi(n)\rangle$ of the $n$ cities such that the total travel distance, calculated as the sum of $d(\pi(i),\pi(i+1))$ for $i=1$ to $n-1$, plus $d(\pi(n),\pi(1))$, is at most $B$?
@@
  • Question: Does $G$ contain a dominating set of size at most $k$, i.e. a subset $W$ of $V$ containing at most $k$ nodes such that every node $u$ in $V-W$ (i.e. in $V$ but not in $W$) has at least one neighbor $w$ in $W$ where ${u,w}$ is an edge in $E$?
@@
  • Question: Can $S$ be partitioned into two disjoint subsets - $S1$ and $S2$ such that for each set $C_{i}$ it holds that $C_{i}$ is not a subset of $S_{1}$ and $C_{i}$ is not a subset of $S_{2}$?
@@
  • Question: Are there $k$ sets - $D_{1},\dots,D_{k}$ from the collection $C$ such that for all $1\leq i<j\leq k$, $D_{i}$ and $D_{j}$ have no common elements?
@@
  • Question: Is there a partition of $X$ into 3 disjoint sets - $X_{1},X_{2},X_{3}$ with which: for each set $X_{i},i\in\{1,2,3\}$, for all pairs $x$ and $y$ in $X_{i}$, it holds that $d(x,y)\leq B$?
@@ Appendix G Full Results over Problems
In this section, we present the full results over problems, as displayed in Figure 8. For each element in the table $x^{a}_{b}$, $x$ is the value of IQM and $a$ and $b$ are the upper and lower values of the CI, respectively.
@@
QwQ-32B $0.94_{0.90}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.56_{0.53}^{0.60}$ $0.11_{0.07}^{0.13}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.83_{0.73}^{0.90}$ $0.52_{0.40}^{0.70}$ $0.32_{0.30}^{0.37}$ $0.19_{0.13}^{0.27}$ $0.13_{0.07}^{0.23}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.84_{0.83}^{0.87}$ $0.27_{0.20}^{0.30}$ $0.17_{0.10}^{0.27}$ $0.08_{0.03}^{0.10}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.94_{0.93}^{0.97}$ $0.51_{0.47}^{0.57}$ $0.43_{0.40}^{0.47}$ $0.22_{0.20}^{0.27}$ $0.09_{0.00}^{0.17}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.89_{0.80}^{0.93}$ $0.62_{0.60}^{0.67}$ $0.54_{0.50}^{0.60}$ $0.36_{0.27}^{0.47}$ $0.19_{0.13}^{0.27}$ $0.14_{0.03}^{0.23}$ $0.08_{0.03}^{0.10}$ $0.03_{0.00}^{0.07}$ $0.02_{0.00}^{0.03}$
DeepSeek-V3 $0.94_{0.93}^{0.97}$ $0.78_{0.60}^{0.90}$ $0.38_{0.33}^{0.40}$ $0.34_{0.17}^{0.43}$ $0.21_{0.17}^{0.27}$ $0.06_{0.03}^{0.10}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.89_{0.83}^{0.97}$ $0.68_{0.60}^{0.80}$ $0.53_{0.47}^{0.63}$ $0.38_{0.30}^{0.43}$ $0.28_{0.23}^{0.33}$ $0.12_{0.03}^{0.23}$ $0.08_{0.03}^{0.17}$ $0.03_{0.03}^{0.03}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.98_{0.93}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.91_{0.87}^{0.97}$ $0.83_{0.63}^{0.93}$ $0.64_{0.63}^{0.67}$ $0.23_{0.20}^{0.27}$ $0.13_{0.10}^{0.17}$
o1-mini $0.92_{0.90}^{0.93}$ $0.91_{0.87}^{0.97}$ $0.92_{0.90}^{0.97}$ $0.81_{0.77}^{0.87}$ $0.67_{0.60}^{0.77}$ $0.20_{0.10}^{0.37}$ $0.03_{0.03}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
o3-mini $0.93_{0.90}^{0.97}$ $0.82_{0.77}^{0.87}$ $0.72_{0.63}^{0.83}$ $0.77_{0.70}^{0.83}$ $0.82_{0.80}^{0.83}$ $0.71_{0.67}^{0.77}$ $0.60_{0.53}^{0.70}$ $0.30_{0.20}^{0.43}$ $0.13_{0.10}^{0.17}$ $0.12_{0.03}^{0.17}$
@@
QwQ-32B $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.50_{0.37}^{0.60}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.91_{0.83}^{1.00}$ $0.92_{0.90}^{0.93}$ $0.81_{0.73}^{0.87}$ $0.52_{0.43}^{0.60}$ $0.03_{0.00}^{0.07}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.94_{0.87}^{1.00}$ $0.67_{0.57}^{0.80}$ $0.37_{0.27}^{0.43}$ $0.18_{0.10}^{0.23}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.96_{0.90}^{1.00}$ $0.88_{0.83}^{0.90}$ $0.78_{0.67}^{0.87}$ $0.60_{0.57}^{0.63}$ $0.01_{0.00}^{0.03}$ $0.03_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.97_{0.90}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.90_{0.90}^{0.90}$ $0.53_{0.47}^{0.57}$ $0.37_{0.30}^{0.47}$ $0.37_{0.23}^{0.50}$ $0.26_{0.20}^{0.30}$ $0.14_{0.10}^{0.17}$ $0.04_{0.00}^{0.07}$
DeepSeek-V3 $0.92_{0.87}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.96_{0.93}^{0.97}$ $0.89_{0.83}^{0.93}$ $0.34_{0.23}^{0.43}$ $0.14_{0.10}^{0.20}$ $0.06_{0.03}^{0.10}$ $0.03_{0.00}^{0.07}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.87_{0.83}^{0.90}$ $0.28_{0.10}^{0.43}$ $0.37_{0.23}^{0.50}$ $0.27_{0.23}^{0.33}$ $0.09_{0.07}^{0.13}$ $0.09_{0.07}^{0.10}$ $0.01_{0.00}^{0.03}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.91_{0.87}^{0.97}$ $0.77_{0.70}^{0.87}$ $0.41_{0.33}^{0.47}$ $0.18_{0.13}^{0.20}$ $0.13_{0.08}^{0.20}$ $0.06_{0.00}^{0.10}$
o1-mini $0.74_{0.73}^{0.77}$ $0.77_{0.73}^{0.80}$ $0.78_{0.70}^{0.83}$ $0.91_{0.87}^{0.93}$ $0.58_{0.43}^{0.70}$ $0.31_{0.27}^{0.33}$ $0.13_{0.10}^{0.17}$ $0.13_{0.03}^{0.27}$ $0.08_{0.07}^{0.10}$ $0.02_{0.00}^{0.07}$
o3-mini $0.82_{0.70}^{0.90}$ $0.89_{0.83}^{0.93}$ $0.89_{0.83}^{0.93}$ $0.80_{0.73}^{0.90}$ $0.59_{0.53}^{0.70}$ $0.52_{0.50}^{0.57}$ $0.19_{0.10}^{0.27}$ $0.13_{0.03}^{0.23}$ $0.11_{0.03}^{0.17}$ $0.07_{0.03}^{0.10}$

### v2

@@ C.1 Problem Suite: npgym
Interface. We introduce npgym, a problem suite containing 25 NPC problems with a unified gym-style interface for instance generation and solution verification. Each environment is defined by a problem name and its corresponding hyperparameters, enabling the generation of unlimited problem instances and example solutions. Difficulty can be scaled by adjusting these parameters. npgym also supports automatic verification of solutions produced by large language models (LLMs). New problems can be added easily by implementing two core functions and providing a problem description for prompt generation.
@@
Variables to Scale. Table 6 lists the variables to scale for each of the 25 NP-complete problems.
@@
Difficulty Levels. We define and release problem-specific difficulty levels for each of the 25 core problems included in our benchmark. Each problem includes approximately 10 levels of increasing complexity, determined primarily by theoretical factors such as search space size and validated through empirical testing using DeepSeek-R1 and GPT-4o. npgym allows seamless extension to higher difficulty levels as more powerful models become available.
@@
Solution Errors. There are two fundamental error categories: problem-independent errors and problem-dependent errors. Problem-independent errors are general errors that arise from external factors unrelated to the problem’s intrinsic characteristics and all problems have these types of errors. Problem-independent errors include JSON ERROR (JSON not found or JSON parsing errors), and VERIFICATION ERROR (output format mismatches or structural validation failures). Problem-dependent errors originate from the problem’s inherent complexity, which are defined based on problem specificity. A comprehensive illustration of the errors is displayed in Table 7.
@@ C.2 Solver Suite: npsolver
Online. The online state-of-the-art LLMs, e.g., o1/o3-mini and DeepSeek-v3/R1, can be accessed through APIs without local computational overhead. However, these online models have dependency on network stability and API costs with token usage. npsolver supports multiple providers, e.g., OpenAI, through modular API clients. We implement efficient batch processing with LiteLLM, which minimizes the latency during parallel problem-solving.
Offline. Open-weight LLMs, e.g., QwQ-32B and Deepseek-R1-32B, can be accessed by deploying them locally. This allows for GPU-accelerated, high-throughput inference while avoiding API-related costs. Offline models are deployed using vLLM, with hyperparameters—such as temperature and maximum token length—manually configured according to their official technical documentation.
@@ C.3 Evaluation Suite: npeval
npeval employs a statistically rigorous sampling strategy. For each difficulty, the aggregated performance over 3 different independent seeds, with 30 samples generated per seed, aligning with the minimum sample size for reliable statistical analysis (Hogg et al., 1977), are considered. This sampling design, i.e., sampling 90 instances total per difficulty level for each problem, balances budget constraints while mitigating instance-specific variance.
Evaluation Metrics. rliable (Agarwal et al., 2021) is an open-source Python library designed to enable statistically robust evaluation of reinforcement learning and machine learning benchmarks. Inspired by rliable, npeval provides the following 4 evaluation aggregate metrics:
@@
To quantify uncertainty in aggregate metrics, e.g. IQM, npeval employs stratified bootstrap confidence intervals (SBCIs) (Efron, 1979; 1987) for the performance interval estimation. SBCIs use stratified resampling within predefined strata, e.g., difficulty levels, to preserve the hierarchical structure of the evaluation data, reduce bias, and provide statistically sound interval estimates.
Comprehensive Analysis. Based on evaluation metrics, npeval provides a comprehensive analysis of the LLMs’ performance over the problems and difficulty levels, including the full results for each problem, each model and each level (Appendix G), the performance over different problems (Appendix H), the analysis of both prompt and completion tokens of LLMs (Appendix I), the analysis of the number of “aha moments” during the DeepSeek-R1 reasoning (Guo et al., 2025) (Appendix J), an illustration of errors over problems (Table 7) with detailed error analysis (Appendix K), considering both the solution errors, i.e., the errors returned by npgym, and the reasoning errors, i.e., the errors produced in the internal reasoning process of LLMs, which enables the identification of the failure cases (Appendix L).
@@ Appendix D Prompts and Responses
This extraction function employs a robust, multi-stage pattern matching approach to handle various response formats. Specifically, this function progressively relaxes format requirements through three regex patterns:
  • First tries to find content between triple quotes with “json” marker,
  • If that fails, looks for “json” followed by content in curly braces,
@@
This flexible parsing template effectively handles format variations produced by different models. For instance, it accommodates models like QwQ-32B that frequently omit format prefixes or deviate from strict formatting instructions. By tolerating minor format violations, this approach ensures that evaluation focuses primarily on models’ reasoning capabilities rather than their adherence to output formatting conventions. Only when all three attempts fail, the JSON error is raised, indicating a genuine failure to produce a feasible solution.
@@ Appendix E List of NP-complete Problems
  • Input: A set of $m$ clauses $\{C_{1},C_{2},\dots,C_{m}\}$ - over a set of $n$ Boolean valued variables $X_{n}=\{x_{1},x_{2},\dots,x_{n}\}$, such that each clause depends on exactly three distinct variables from $X_{n}$. A clause being a Boolean expression of the form $y_{i}\land y_{j}\land y_{k}$ where each $y$ is of the form $x$ or $\neg x$ (i.e. negation of $x$) with $x$ being some variable in $X_{n}$. For example if $n=4$ and $m=3$, a possible instance could be the (set of) Boolean expressions: $C_{1}=(x_{1}\land(\neg x_{2})\land(\neg x_{3}))$, $C_{2}=(x_{2}\land x_{3}\land(\neg x_{4}))$, $C_{3}=((\neg x_{1})\land x_{3}\land x_{4})$.
@@
  • Question: Are there two positive integers $x$ and $y$ such that (axx)+(by)=c(a*x*x)+(b*y)=c?
@@
  • Question: Is there a linear ordering of $V$ with bandwidth at most $k$, i.e., a one-to-one function $f:V\rightarrow\{0,1,2,...,n-1\}$ such that for all edges u,v{u,v} in $G$, $|f(u)-f(v)|\leq k$?
@@
  • Input: A set $C$ of n cities $\{c_{1},\dots,c_{n}\}$; a positive integer distance $d(i,j)$ for each pair of cities $(c_{i},c_{j}),i<j,i,j\in\{1,\dots,n\}$; a positive integerBB representing the maximum allowed travel distance.
  • Question: Is there an ordering $\langle\pi(1),\pi(2),...,\pi(n)\rangle$ of the $n$ cities such that the total travel distance, calculated as the sum of $d(\pi(i),\pi(i+1))$ for $i=1$ to n1n-1, plus $d(\pi(n),\pi(1))$, is at most $B$?
@@
  • Question: Does $G$ contain a dominating set of size at most $k$, i.e. a subset $W$ of $V$ containing at most $k$ nodes such that every node $u$ in VWV-W (i.e. in $V$ but not in $W$) has at least one neighbor $w$ in $W$ where u,w{u,w} is an edge in $E$?
@@
  • Question: Can $S$ be partitioned into two disjoint subsets - $S1$ and $S2$ - such that for each set $C_{i}$ it holds that $C_{i}$ is not a subset of $S_{1}$ and $C_{i}$ is not a subset of $S_{2}$?
@@
  • Question: Are there $k$ sets - $D_{1},\dots,D_{k}$ - from the collection $C$ such that for all $1\leq i<j\leq k$, $D_{i}$ and $D_{j}$ have no common elements?
@@
  • Question: Is there a partition of $X$ into 3 disjoint sets - $X_{1},X_{2},X_{3}$ - with which: for each set $X_{i},i\in\{1,2,3\}$, for all pairs $x$ and $y$ in $X_{i}$, it holds that $d(x,y)\leq B$?
@@ Appendix G Full Results over Problems
In this section, we present the full results over problems, as displayed in Figure 6. For each element in the table $^{a}_{b}, xx$ is the value of IQM and $a$ and $b$ are the upper and lower values of the CI, respectively.
@@
QwQ-32B $0.94_{0.90}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.56_{0.53}^{0.60}$ $11_{0.07}^{0.13} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
DeepSeek-R1-32B $0.83_{0.73}^{0.90}$ $0.52_{0.40}^{0.70}$ $0.32_{0.30}^{0.37}$ $0.19_{0.13}^{0.27}$ $13_{0.07}^{0.23} 0.0$$2_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o-mini $0.84_{0.83}^{0.87}$ $0.27_{0.20}^{0.30}$ $17_{0.10}^{0.27} 0.0$$8_{0.03}^{0.10} 0.0$$2_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
GPT-4o $0.94_{0.93}^{0.97}$ $0.51_{0.47}^{0.57}$ $0.43_{0.40}^{0.47}$ $22_{0.20}^{0.27} 0.0$$9_{0.00}^{0.17} 0.0$$1_{0.00}^{0.03} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.89_{0.80}^{0.93}$ $0.62_{0.60}^{0.67}$ $0.54_{0.50}^{0.60}$ $0.36_{0.27}^{0.47}$ $0.19_{0.13}^{0.27}$ $14_{0.03}^{0.23} 0.0$$8_{0.03}^{0.10} 0.0$$3_{0.00}^{0.07} 0.0$$02_{0.00}^{0.03}$
DeepSeek-V3 $0.94_{0.93}^{0.97}$ $0.78_{0.60}^{0.90}$ $0.38_{0.33}^{0.40}$ $0.34_{0.17}^{0.43}$ $21_{0.17}^{0.27} 0.0$$6_{0.03}^{0.10} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.89_{0.83}^{0.97}$ $0.68_{0.60}^{0.80}$ $0.53_{0.47}^{0.63}$ $0.38_{0.30}^{0.43}$ $0.28_{0.23}^{0.33}$ $12_{0.03}^{0.23} 0.0$$8_{0.03}^{0.17} 0.0$$03_{0.03}^{0.03}$
DeepSeek-R1 $_{1.00}^{1.00} 1.001.00$$_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.98_{0.93}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.91_{0.87}^{0.97}$ $0.83_{0.63}^{0.93}$ $0.64_{0.63}^{0.67}$ $0.23_{0.20}^{0.27}$ $0.13_{0.10}^{0.17}$
o1-mini $0.92_{0.90}^{0.93}$ $0.91_{0.87}^{0.97}$ $0.92_{0.90}^{0.97}$ $0.81_{0.77}^{0.87}$ $0.67_{0.60}^{0.77}$ $20_{0.10}^{0.37} 0.0$$3_{0.03}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
o3-mini $0.93_{0.90}^{0.97}$ $0.82_{0.77}^{0.87}$ $2_{0.63}^{0.83} 0.770$$7_{0.70}^{0.83}$ $0.82_{0.80}^{0.83}$ $0.71_{0.67}^{0.77}$ $0.60_{0.53}^{0.70}$ $0.30_{0.20}^{0.43}$ $0.13_{0.10}^{0.17}$ $0.12_{0.03}^{0.17}$
@@
QwQ-32B $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.93_{0.90}^{0.97}$ $50_{0.37}^{0.60} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
DeepSeek-R1-32B $0.91_{0.83}^{1.00}$ $0.92_{0.90}^{0.93}$ $0.81_{0.73}^{0.87}$ $52_{0.43}^{0.60} 0.0$$3_{0.00}^{0.07} 0.0$$2_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o-mini $0.94_{0.87}^{1.00}$ $0.67_{0.57}^{0.80}$ $0.37_{0.27}^{0.43}$ $18_{0.10}^{0.23} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o $0.96_{0.90}^{1.00}$ $0.88_{0.83}^{0.90}$ $0.78_{0.67}^{0.87}$ $60_{0.57}^{0.63} 0.0$$1_{0.00}^{0.03} 0.0$$3_{0.00}^{0.07} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.97_{0.90}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.90_{0.90}^{0.90}$ $0.53_{0.47}^{0.57}$ $0.37_{0.30}^{0.47}$ $0.37_{0.23}^{0.50}$ $0.26_{0.20}^{0.30}$ $14_{0.10}^{0.17} 0.0$$04_{0.00}^{0.07}$
DeepSeek-V3 $0.92_{0.87}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.96_{0.93}^{0.97}$ $0.89_{0.83}^{0.93}$ $0.34_{0.23}^{0.43}$ $14_{0.10}^{0.20} 0.0$$6_{0.03}^{0.10} 0.0$$3_{0.00}^{0.07} 0.0$$3_{0.00}^{0.07} 0.0$$01_{0.00}^{0.03}$
DeepSeek-V3-2503 $_{1.00}^{1.00} 1.001.00$$_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.87_{0.83}^{0.90}$ $0.28_{0.10}^{0.43}$ $0.37_{0.23}^{0.50}$ $27_{0.23}^{0.33} 0.0$$9_{0.07}^{0.13} 0.0$$9_{0.07}^{0.10} 0.0$$01_{0.00}^{0.03}$
DeepSeek-R1 $_{1.00}^{1.00} 1.001.00$$_{1.00}^{1.00}$ $_{1.00}^{1.00} 1.001.00$$_{1.00}^{1.00}$ $0.91_{0.87}^{0.97}$ $0.77_{0.70}^{0.87}$ $0.41_{0.33}^{0.47}$ $0.18_{0.13}^{0.20}$ $13_{0.08}^{0.20} 0.0$$06_{0.00}^{0.10}$
o1-mini $4_{0.73}^{0.77} 0.770$$7_{0.73}^{0.80}$ $0.78_{0.70}^{0.83}$ $0.91_{0.87}^{0.93}$ $0.58_{0.43}^{0.70}$ $0.31_{0.27}^{0.33}$ $0.13_{0.10}^{0.17}$ $13_{0.03}^{0.27} 0.0$$8_{0.07}^{0.10} 0.0$$02_{0.00}^{0.07}$
o3-mini $0.82_{0.70}^{0.90}$ $0.89_{0.83}^{0.93}$ $0.89_{0.83}^{0.93}$ $0.80_{0.73}^{0.90}$ $0.59_{0.53}^{0.70}$ $0.52_{0.50}^{0.57}$ $0.19_{0.10}^{0.27}$ $3_{0.03}^{0.23} 0.110$$1_{0.03}^{0.17}$ $0.07_{0.03}^{0.10}$